As we know the thermal decomposition of ethyl bromide is complex, the overall
rate is first-order, and the lab constant is given by the expression k = (3.8 x
10^14 s-1)e-229000/RT, where the activation energy is in J mol-1. Calculate the
temperature at which (a) ethyl bromide decomposes at the rate of 1% per
second and (b) the decomposition is 70% complete in 1 hour.
1
Expert's answer
2020-10-21T13:18:42-0400

For the first order reaction, "[X]=[X]_ie^{-kt}"

"(a)" "[X]=\frac{99}{100}[X]_i"

"ln(\frac{99}{100})=-kt" at "t=1 sec"

"-0.01005=(-3.8\times10^{14})e^{-\frac{229000}{RT}}"

Taking log both sides;

"-4.600=1.335+32.236-\frac{229000}{RT}"

"\frac{229000}{8.314\times T}=38.171"

Hence, "T=721.588K"


"(b)" "ln(0.3)=-(3.8\times10^{14})e^{-\frac{229000}{RT}}\times3600"

"-1.2039=-1368\times10^{15}\times e^{-\frac{229000}{RT}}"

"8.8\times10^{-19}=e^{-\frac{229000}{RT}}"

"16.825=\frac{229000}{8.314\times T}"

Hence, "T=1637.05K"